1. A 600 MW steam power plant, which is cooled by a river, has a thermal efficiency of 40 percent. Determine the rate of heat transfer to the river water. Will the actual heat transfer rate be higher or lower than this value? Why?

1
Expert's answer
2022-02-25T09:07:02-0500

"efficiency=\frac{W_{net}}{rate\space of \space heat \space transfer}"rate of heat transfer "=\frac{W_{net}}{efficiency}=\frac{600}{40/100}=1500MW"

The rate of heat transfer to the river water is determined from the first law relation for a heat engine.

"\dot{Q}_L=\dot{Q}_H-\dot{W}_{net out}=1500-600=900MW"

The amount of heat rejected to the river will be lower since part of the heat will be lost to the surrounding air from the working fluid as it passes through the pipes and other components.